# Lieb-Schultz-Mattis Anomalies and Anomaly Matching

### 01 Apr 2026 v1

@@ 3 LSM IN HIGHER DIMENSIONS
Many of the results discussed in Sec. 2 can be generalized to higher dimensions, which we review below. Before diving in, it is worth highlighting some important differences between higher dimensional systems and 1D systems. First, the crystalline symmetries are much richer in higher dimensions. There are only 2 space groups in 1D, but there are 17 space groups in 2D and 230 in 3D. Second, as discussed in Sec. 2, a 1D system with a LSM anomaly may spontaneously break the anomalous symmetry or be gapless. In higher dimensions, new possibilities arise, i.e., it can also be a symmetric gapped topological phase. In this section, we focus on the LSM constraints themselves. In Sec. 4, we review the general framework for anomaly matching and demonstrate it using various higher dimensional examples.
@@
With the symmetry setting above, if each unit cell contains a spin-1/2 moment and if the interactions in the Hamiltonian decays as $1/r^{a}$ with a>4d2a>4d-2 ($r$ is the range of the interaction), the Hamiltonian cannot have a unique gapped ground state.
@@
In other symmetry settings, where the on-site symmetry may be discrete and the crystalline symmetry may go beyond translation, rigorous results are rare, but an important picture known as “lattice homotopy” is very powerful [4, 8]. In essence, given a symmetry group, lattice homotopy is a way to organize the infinitely many types of lattice systems to a finite set of equivalence classes, such that two lattice systems have different (the same) anomalies if they are in different (the same) classes. The simplest setup to demonstrate the notion of lattice homotopy is lattice systems with symmetry $G_{\rm s}\times G_{\rm int}$, where $G_{\rm s}$ is a space group symmetry and $G_{\rm int}$ is an on-site symmetry. Then two lattice systems are in the same lattice homotopy class if and only if they can be deformed into each other by these operations: 1) moving the microscopic DOF while preserving $G_{\rm s}$, 2) at each location, identifying DOF with the same type of projective representation under $G_{\rm int}$, in a way preserving $G_{\rm s}$.
@@
More generally, one can see that all lattice systems with $p6m\times\mathrm{SO}(3)$ symmetry are classified into 4 types by lattice homotopy, with representative systems being a lattice with only spin-singlet DOF, a triangular lattice spin-1/2 system, a kagome lattice spin-1/2 system, and a system with spin-1/2 moments at both the triangular and kagome sites. The first among the four is anomaly free, and the other three all have different LSM anomalies. The anomaly-free systems allow symmetric SRE state and symmetric local gapped Hamiltonians with a unique ground state, while the anomalous systems do not. We remark that the last among the four types has an even number of spin-1/2 moments per unit cell, so with only translation and $\mathrm{SO}(3)$ symmetries it is anomaly free, and the anomaly is actually a mixed anomaly between the $\mathrm{SO}(3)$ and lattice $C_{2}$ rotation symmetries.
Figure 1: Panel (a) shows the generators of the $p6m$ group, including translations $T_{1}$ and $T_{2}$, a 6-fold rotation $C_{6}$ and a mirror reflection $M$. The two translation vectors have the same length, and their angle is $2\pi/3$. The reflection axis of $M$ bisects these two translation vectors. In panel (b), the hexagon is a translation unit cell of the $p6m$ lattice symmetry. There are three irreducible Wyckoff positions, labelled by $a$, $b$ and $c$, and they form the sites of the triangular, honeycomb and kagome lattices, respectively. The $C_{6}$ rotation center in panel (a) is at a type-$a$ point.
@@ 4.1 General Theory of UV-IR Matching
An IR theory can be characterized by its symmetries. We are mostly interested in its internal symmetries, as the (continuous) spacetime symmetries in the continuum theories do not play much role in anomaly matching. Notably, it is quite common that the IR theory embodies internal symmetries not present in the UV. We have seen such an example in spin-1/2 Heisenberg chain, where the deep IR theory (the SU(2)1 CFT) has SO(4) symmetry, while the UV lattice model only has SO(3) internal symmetry. Symmetries that are present in the IR theory but absent in UV are called “emergent”. From the perspective of renormalization group (RG), emergent symmetries are broken by IR irrelevant perturbations.
Denote the symmetry of the UV system by $G_{\rm UV}$. Suppose the IR theory has (emergent) symmetry $G_{\rm IR}$. Since the UV and IR theories are connected by the RG flow, all UV symmetries must have their counterparts in IR [^4] It is possible that a UV symmetry acts trivially in the IR theory., which, mathematically, are captured by a group homomorphism $\varphi$ between them:
@@
This mapping $\varphi$ describes the “symmetry enrichment” of the IR theory. Note that given the same $G_{\rm UV}$ and $G_{\rm IR}$ (and the same IR theory), there can be multiple choices of $\varphi$’s, which then correspond to physically distinct “symmetry-enriched” phases.
Besides the group homomorphism, to examine whether an IR theory is valid for a given UV system, we also need to account for the ’t Hooft anomaly, which roughly speaking describes certain topological response of the system to defects, as reviewed in Sec. 2.1 for the LSM anomaly in spin chains. The anomaly is independent of energy scale and remains invariant under RG flows, so the IR theory must produce the same response, which was demonstrated in a concrete example in Sec. 2.3. Here we will present a more general and compact formalism.
@@
In order to discuss the anomaly matching condition, we assume that the anomaly is abstractly described by a “function” $\omega_{\rm UV}$ of $A_{\rm UV}$, which gives the response of the system when the background field $A_{\rm UV}$ is turned on [^5] If the system is in continuum and Lorentz-invariant, then $\omega$ is more conveniently presented as a response action $S_{\rm UV}[A_{\rm UV}]$ in one-dimensional higher.. The two anomalies $\omega_{\rm UV}[A_{\rm UV}]$ and $\omega_{\rm IR}[A_{\rm IR}]$ are then related as:
@@
    So far, all symmetries under consideration act on the entire system. In other words, charged objects are point-like objects, i.e., particles. However, this is not the full story in many IR theories. The need to extend the notion of symmetries is best illustrated when the IR theory is a topological quantum field theory (TQFT). In general, TQFTs can only have finite global symmetry that acts faithfully (in the sense defined so far), and naively the matching condition cannot be satisfied. The resolution is that one has to consider a more general notion of symmetries, whose charged objects are not point-like, but are extended in spacetime (e.g., loops in 2+1d). These “generalized” symmetries are called higher-form symmetries [55]. However, in Sec. 4.4, we will give another approach for anomaly matching in a topological IR theory without using the language of higher-form symmetries.
@@ 4.2 LSM Anomalies for Various Lattice Systems
The UV anomaly, in its simplest case that only involves the lattice translation symmetry and another on-site symmetry, was derived in Ref. [27]. The generalizations to include an arbitrary space group symmetry were made in Ref. [28] for 1D and 2D systems and in Ref. [31] for 3D systems. Concretely, for any lattice system in $d=1,2,3$ spatial dimension with symmetry group $G_{\rm UV}=G_{\rm s}\times G_{\rm int}$, where $G_{\rm s}$ is an arbitrary space group and $G_{\rm int}$ is an on-site symmetry whose projective representations are classified by $\mathbb{Z}_{2}^{k}$ (i.e., there exists a positive integer $k$ such that $H^{2}(G_{\rm int},\mathrm{U}(1))=\mathbb{Z}_{2}^{k}$), then $\omega_{\rm UV}$ can be expressed as an element in the group cohomology $H^{d+1}(G_{\rm UV},\mathrm{U}(1))$:
@@
A few remarks are in order. First, the symmetry settings above cover a large class of lattice systems relevant to experimental and numerical studies, especially in the context of quantum spin liquids, where $G_{\rm int}$ can be the $\mathrm{SO}(3)$ spin rotation, $\mathbb{Z}_{2}^{\mathsf{T}}$ time reversal, etc. Second, we note that Eq. (19) takes a notably simple factorized form, where the information of the type of lattice is completely encoded in $\lambda$, and the information of the projective representation of the DOF is completely encoded in $\eta$. Moreover, given the type of lattice, $\lambda$ has a universal expression for all $G_{\rm int}$ and all $\eta$. We also note that not all elements in $H^{d-1}(G_{\rm s},\mathbb{Z}_{2})$ can be a valid $\lambda$, and an important contribution of Refs. [28, 31] is to determine the $\lambda$ for each lattice.
@@
Now we can ask how the $p6m$ symmetry acts on the $\mathrm{SO}(3)$ monopole, and the different symmetry actions are precisely classified by $H^{2}(p6m,\mathbb{Z}_{2})$ [56]. For example, although the two translations $T_{1,2}$ in Figure 1 should commute on local operators, their actions on the $SO(3)$ monopole may anti-commute, signaling a ’t Hooft anomaly associated with the translation and $\mathrm{SO}(3)$ symmetries. Now consider the effect of $T_{2}^{-1}T_{1}^{-1}T_{2}T_{1}$ on the $SO(3)$ monopole. This operation moves this monopole around a unit cell. Suppose the unit cell includes a half-integer (integer) spin, this process is analogous to moving a $2\pi$ magnetic flux around a half-integer (integer) charge, so a 1-1 (1) Aharonov-Bohm phase factor will be generated and $T_{1,2}$ anti-commute (commute) on the monopole. So the spin moments in a unit cell dictate the mixed anomaly between the translation and $SO(3)$ symmetries. Continuing this line of reasoning, one can see that the type of the lattice, or more precisely, the position and the spin quantum number of the magnetic moments, completely determines how the entire $p6m$ symmetry acts on an $\mathrm{SO}(3)$ monopole and therefore fixes an element in $H^{2}(p6m,\mathbb{Z}_{2})$. This element is precisely $\lambda$. On the other hand, the spin quantum numbers determine $\eta$. Altogether, $\lambda$ and $\eta$ determine the UV anomaly via Eq. (19).
@@ 4.3 Dirac Spin Liquids
The first example is the Dirac spin liquid (DSL), which was originally proposed in the context of high-temperature superconductors [57, 58]. Recently, evidence of the emergence of DSL has been reported in kagome and triangular anti-ferromagnets [59, 60, 61, 62, 63, 64, 65, 66, 67]. An intuitive way to think of the DSL is to “fractionalize” each localized spin into two fermions, and these fermions experience some band structure that has 4 Dirac cones (including all spins and valleys). As common in the phenomena of fractionalization, these fermions are coupled to an emergent dynamical gauge field, which, in this case, is a $\mathrm{U}(1)$ gauge field. At low energies, there is evidence that the DSL is described by an emergent CFT [68, 69, 70].
It is worth noting that it is more appropriate to view the fundamental low-energy DOF of the DSL as the Dirac monopoles, rather than the Dirac fermions, because the latter is not gauge invariant, while the former is, and from the former (and its conjugate variable, which is the field strength of the $\mathrm{U}(1)$ gauge field) all gauge invariant operators can be constructed. The monopole operators in this DSL can be organized into a 6-by-2 matrix field $n$ with Hermitian entries [71].
@@
$\begin{split}&T_{1}:n\rightarrow\left(\begin{array}[]{cccc}I_{3}&&&\\ &1&&\\ &&-1&\\ &&&-1\\ \end{array}\right)n\left(\begin{array}[]{cc}-\frac{1}{2}&-\frac{\sqrt{3}}{2}\\ \frac{\sqrt{3}}{2}&-\frac{1}{2}\end{array}\right),\quad T_{2}:n\rightarrow\left(\begin{array}[]{cccc}I_{3}&&&\\ &-1&&\\ &&1&\\ &&&-1\\ \end{array}\right)n\left(\begin{array}[]{cc}-\frac{1}{2}&-\frac{\sqrt{3}}{2}\\ \frac{\sqrt{3}}{2}&-\frac{1}{2}\end{array}\right),\\ &C_{6}:n\rightarrow\left(\begin{array}[]{cccc}I_{3}&&&\\ &&1&\\ &&&1\\ &-1&&\\ \end{array}\right)n\left(\begin{array}[]{cc}1&\\ &-1\\ \end{array}\right),\quad M:n\rightarrow\left(\begin{array}[]{cccc}I_{3}&&&\\ &&-1&\\ &-1&&\\ &&&1\\ \end{array}\right)n,\quad{\rm O}(3)^{\mathsf{T}}:n\rightarrow\left(\begin{array}[]{cc}{\rm O}(3)^{\mathsf{T}}&\\ &I_{3}\\ \end{array}\right)n,\end{split}$ (21)
@@
Besides the above symmetry-enriched pattern that had been discovered, the anomaly-based classification in Ref. [28] predicted various exotic symmetry-enrichment patterns of DSL unknown before, as well as symmetry-enrichment patterns of some even more exotic “non-Lagrangian” quantum matter [71]. It still remains to find concrete models to realize these predicted phases.

### 29 Apr 2026 v2

@@ 3 LSM IN HIGHER DIMENSIONS
Many of the results discussed in Sec. 2 can be generalized to higher dimensions, which we review below. Before diving in, it is worth highlighting some important differences between higher-dimensional systems and 1D systems. First, the crystalline symmetries are much richer in higher dimensions. There are only 2 space groups in 1D, but there are 17 space groups in 2D and 230 in 3D. Second, as discussed in Sec. 2, a 1D system with a LSM anomaly may spontaneously break the anomalous symmetry or be gapless. In higher dimensions, new possibilities arise, e.g., it can also be a symmetric gapped topological phase. In this section, we focus on the LSM constraints themselves. In Sec. 4, we review the general framework for anomaly matching and demonstrate it using various higher-dimensional examples.
@@
With the symmetry setting above, if each unit cell contains a spin-1/2 moment and if the interactions in the Hamiltonian decay as $1/r^{a}$ with a>4d2a>4d-2 ($r$ is the distance between the interacting spins), the Hamiltonian cannot have a unique gapped ground state.
@@
In other symmetry settings, where the on-site symmetry may be discrete and the crystalline symmetry may go beyond translation, rigorous results are rare, but an important picture known as “lattice homotopy” is very powerful [4, 8]. In essence, given a symmetry group, lattice homotopy is a way to organize the infinitely many types of lattice systems into a finite set of equivalence classes, such that two lattice systems have different (the same) anomalies if they are in different (the same) classes. The simplest setup to demonstrate the notion of lattice homotopy is lattice systems with symmetry $G_{\rm s}\times G_{\rm int}$, where $G_{\rm s}$ is a space group symmetry and $G_{\rm int}$ is an on-site symmetry. Then two lattice systems are in the same lattice homotopy class if and only if they can be deformed into each other by these operations: 1) moving the microscopic DOF while preserving $G_{\rm s}$, 2) at each location, identifying DOF with the same type of projective representation under $G_{\rm int}$, in a way preserving $G_{\rm s}$.
@@
More generally, one can see that all lattice systems with $p6m\times\mathrm{SO}(3)$ symmetry are classified into 4 types by lattice homotopy, with representative systems being a lattice with only spin-singlet DOF, a triangular lattice spin-1/2 system, a kagome lattice spin-1/2 system, and a system with spin-1/2 moments at both the triangular and kagome sites. The first among the four is anomaly free, and the other three all have different LSM anomalies. The anomaly-free systems allow a symmetric SRE state and symmetric local gapped Hamiltonians with a unique ground state, while the anomalous systems do not. We remark that the last among the four types has an even number of spin-1/2 moments per unit cell, so with only translation and $\mathrm{SO}(3)$ symmetries it is anomaly free, and the anomaly is actually a mixed anomaly between the $\mathrm{SO}(3)$ and lattice $C_{2}$ rotation symmetries.
@@ 4.1 General Theory of UV-IR Matching
An IR theory can be characterized by its symmetries. We are mostly interested in its internal symmetries, as the (continuous) spacetime symmetries in the continuum theories do not play much role in anomaly matching. Notably, it is quite common that the IR theory embodies internal symmetries not present in the UV. We have seen such an example in the spin-1/2 Heisenberg chain, where the deep IR theory (the SU(2)1 CFT) has SO(4) symmetry, while the UV lattice model only has SO(3) internal symmetry. Symmetries that are present in the IR theory but absent in the UV are called “emergent”. From the perspective of renormalization group (RG), emergent symmetries are broken by IR irrelevant perturbations.
Denote the symmetry of the UV system by $G_{\rm UV}$. Suppose the IR theory has (emergent) symmetry $G_{\rm IR}$. Since the UV and IR theories are connected by the RG flow, all UV symmetries must have their counterparts in the IR [^3] It is possible that a UV symmetry acts trivially in the IR theory., which, mathematically, are captured by a group homomorphism $\varphi$ between them:
@@
This mapping $\varphi$ describes the “symmetry enrichment” of the IR theory. Note that given the same $G_{\rm UV}$ and $G_{\rm IR}$ (and the same IR theory), there can be multiple choices of $\varphi$, which then correspond to physically distinct “symmetry-enriched” phases.
Besides the group homomorphism, to examine whether an IR theory is valid for a given UV system, we also need to account for the ’t Hooft anomaly, which describes, roughly speaking, certain topological responses of the system to defects, as reviewed in Sec. 2.1 for the LSM anomaly in spin chains. The anomaly is independent of energy scale and remains invariant under RG flows, so the IR theory must produce the same response, which was demonstrated in a concrete example in Sec. 2.3. Here we will present a more general and compact formalism.
@@
In order to discuss the anomaly matching condition, we assume that the anomaly is abstractly described by a “function” $\omega_{\rm UV}$ of $A_{\rm UV}$, which gives the response of the system when the background field $A_{\rm UV}$ is turned on [^4] If the system is in the continuum and Lorentz-invariant, then $\omega$ is more conveniently presented as a response action $S_{\rm UV}[A_{\rm UV}]$ in one dimension higher.. The two anomalies $\omega_{\rm UV}[A_{\rm UV}]$ and $\omega_{\rm IR}[A_{\rm IR}]$ are then related as:
@@
A simple but nontrivial observation from Eq. (18) is that in any microscopic system with a nontrivial anomaly, i.e., $\omega_{\rm UV}\neq 0$, the low-energy theory cannot be a conventional Wilson-Fisher theory, because such Wilson-Fisher theories allow a trivially disordered phase and thus have $\omega_{\rm IR}=0$, which cannot satisfy Eq. (18). However, some “fractionalized” analog of Wilson-Fisher theories can still match the LSM anomalies, and they were indeed identified in various anomalous systems [57, 58].
@@
    So far, all symmetries under consideration act on the entire system. In other words, charged objects are point-like objects, i.e., particles. However, this is not the whole picture in many IR theories. The need to extend the notion of symmetries is best illustrated when the IR theory is a topological quantum field theory (TQFT). In general, TQFTs can only have finite global symmetry that acts faithfully (in the sense defined so far), and naively the matching condition cannot be satisfied. The resolution is that one has to consider a more general notion of symmetries, whose charged objects are not point-like, but are extended in spacetime (e.g., loops in 2+1d). These “generalized” symmetries are called higher-form symmetries [60]. However, in Sec. 4.4, we will give another approach for anomaly matching in a topological IR theory without using the language of higher-form symmetries.
@@ 4.2 LSM Anomalies for Various Lattice Systems
The UV anomaly, in its simplest case that only involves the lattice translation symmetry and another on-site symmetry, was derived in Ref. [27]. The generalizations to include an arbitrary space group symmetry were made in Ref. [31] for 1D and 2D systems and in Ref. [34] for 3D systems. Concretely, for any lattice system in $d=1,2,3$ spatial dimensions with symmetry group $G_{\rm UV}=G_{\rm s}\times G_{\rm int}$, where $G_{\rm s}$ is an arbitrary space group and $G_{\rm int}$ is an on-site symmetry whose projective representations are classified by $\mathbb{Z}_{2}^{k}$ (i.e., there exists a positive integer $k$ such that $H^{2}(G_{\rm int},\mathrm{U}(1))=\mathbb{Z}_{2}^{k}$), then $\omega_{\rm UV}$ can be expressed as an element in the group cohomology $H^{d+1}(G_{\rm UV},\mathrm{U}(1))$:
@@
A few remarks are in order. First, the symmetry settings above cover a large class of lattice systems relevant to experimental and numerical studies, especially in the context of quantum spin liquids, where $G_{\rm int}$ can be the $\mathrm{SO}(3)$ spin rotation, $\mathbb{Z}_{2}^{\mathsf{T}}$ time reversal, etc. Second, we note that Eq. (19) takes a notably simple factorized form, where the information about the type of lattice is completely encoded in $\lambda$, and the information about the projective representation of the DOF is completely encoded in $\eta$. Moreover, given the type of lattice, $\lambda$ has a universal expression for all $G_{\rm int}$ and all $\eta$. We also note that not all elements in $H^{d-1}(G_{\rm s},\mathbb{Z}_{2})$ can be a valid $\lambda$, and an important contribution of Refs. [31, 34] is to determine the $\lambda$ for each lattice.
Figure 1: Panel (a) shows the generators of the $p6m$ group, including translations $T_{1}$ and $T_{2}$, a 6-fold rotation $C_{6}$ and a mirror reflection $M$. The two translation vectors have the same length, and their angle is $2\pi/3$. The reflection axis of $M$ bisects these two translation vectors and its angle with both translation vectors is $\pi/3$. In panel (b), the hexagon is a translation unit cell of the $p6m$ lattice symmetry. There are three irreducible Wyckoff positions, labelled by $a$, $b$ and $c$, and they form the sites of the triangular, honeycomb and kagome lattices, respectively. The $C_{6}$ rotation center in panel (a) is at a type-$a$ point.
@@
Now we can ask how the $p6m$ symmetry acts on the $\mathrm{SO}(3)$ monopole, and the different symmetry actions are precisely classified by $H^{2}(p6m,\mathbb{Z}_{2})$ [61]. For example, although the two translations $T_{1,2}$ in Figure 1 should commute on local operators, their actions on the $SO(3)$ monopole may anti-commute, signaling a ’t Hooft anomaly associated with the translation and $\mathrm{SO}(3)$ symmetries. To see this, consider the effect of $T_{2}^{-1}T_{1}^{-1}T_{2}T_{1}$ on the $SO(3)$ monopole. This operation moves this monopole around a unit cell. Suppose the unit cell includes a half-integer (integer) spin. This process is analogous to moving a $2\pi$ magnetic flux around a half-integer (integer) charge, so a 1-1 (1) Aharonov-Bohm phase factor will be generated and $T_{1,2}$ anti-commute (commute) on the monopole. So the spin moments in a unit cell dictate the mixed anomaly between the translation and $SO(3)$ symmetries. Continuing this line of reasoning, one can see that the type of the lattice, or more precisely, the position and the spin quantum number of the magnetic moments, completely determines how the entire $p6m$ symmetry acts on an $\mathrm{SO}(3)$ monopole and therefore fixes an element in $H^{2}(p6m,\mathbb{Z}_{2})$. This element is precisely $\lambda$. On the other hand, the spin quantum numbers determine $\eta$. Altogether, $\lambda$ and $\eta$ determine the UV anomaly via Eq. (19).
@@ 4.3 Dirac Spin Liquids
The first example is the Dirac spin liquid (DSL), which was originally proposed in the context of high-temperature superconductors [62, 63]. Recently, evidence of the emergence of DSL has been reported in kagome and triangular antiferromagnets [64, 65, 66, 67, 68, 69, 70, 71, 72]. An intuitive way to think of the DSL is to “fractionalize” each localized spin into two fermions, and these fermions experience some band structure that has 4 Dirac cones (including all spins and valleys). As is common in fractionalization phenomena, these fermions are coupled to an emergent dynamical gauge field, which, in this case, is a $\mathrm{U}(1)$ gauge field. At low energies, there is evidence that the DSL is described by an emergent CFT [73, 74, 75].
For our purpose, it is more appropriate to use the gauge-invariant monopole operators as the fundamental low-energy DOF of the DSL, instead of the Dirac fermions because the latter is not gauge invariant. From the monopole operator (and its conjugate variable, which is the field strength of the $\mathrm{U}(1)$ gauge field) all gauge invariant operators can be constructed. The monopole operators in this DSL can be organized into a 6-by-2 matrix field $n$ with Hermitian entries [76].
@@
$\begin{split}&T_{1}:n\rightarrow\left(\begin{array}[]{cccc}I_{3}&&&\\ &1&&\\ &&-1&\\ &&&-1\\ \end{array}\right)n\left(\begin{array}[]{cc}-\frac{1}{2}&-\frac{\sqrt{3}}{2}\\ \frac{\sqrt{3}}{2}&-\frac{1}{2}\end{array}\right),\quad T_{2}:n\rightarrow\left(\begin{array}[]{cccc}I_{3}&&&\\ &-1&&\\ &&1&\\ &&&-1\\ \end{array}\right)n\left(\begin{array}[]{cc}-\frac{1}{2}&-\frac{\sqrt{3}}{2}\\ \frac{\sqrt{3}}{2}&-\frac{1}{2}\end{array}\right),\\ &C_{6}:n\rightarrow\left(\begin{array}[]{cccc}I_{3}&&&\\ &&1&\\ &&&1\\ &-1&&\\ \end{array}\right)n\left(\begin{array}[]{cc}1&\\ &-1\\ \end{array}\right),\quad M:n\rightarrow\left(\begin{array}[]{cccc}I_{3}&&&\\ &&-1&\\ &-1&&\\ &&&1\\ \end{array}\right)n,\\ &{\rm O}(3)^{\mathsf{T}}:n\rightarrow\left(\begin{array}[]{cc}{\rm O}(3)^{\mathsf{T}}&\\ I_{3}\\ \end{array}\right)n,\end{split}$ (21)
@@
Besides the above symmetry-enrichment pattern that has been previously discovered, the anomaly-based classification in Ref. [31] predicted various exotic symmetry-enrichment patterns of DSL that were previously unknown, as well as symmetry-enrichment patterns of some even more exotic “non-Lagrangian” quantum matter [76]. It still remains to find concrete models to realize these predicted phases.